karenfrank:

Karen bakes cookies. She has observed that the number of cookies that turn out well in each batch varies with the size of the batch and is given by the function f(x) = 2x – 23, where x is the size of the batch (number of cookies in each batch). She never bakes fewer than 12 cookies in a batch. If the size of five batches is {20, 25, 26, 27, 34}, how many cookies turn out well in each batch?

HuskyNation:

Just plug each number in and find the answer. f(x) = 2(20) - 23 = 17 f(x) = 2(25) - 23 = 27 f(x) = 2(26) - 23 = 29 f(x) = 2(27) - 23 = 31 f(x) = 2(34) - 23 = 45

HuskyNation:

So the answer is {17, 27, 29, 31, 45}

HuskyNation:

\[\style{fontFamily;Comic Sans;text-shadow: 10px 10px 20px black, 0 0 10px #0000ff, 0 0 5px darkblue;color:black;font-size:100px;font-weight:bold;padding:50px;width:500px;height:75%;position:absolute;left:0;z-index:500;text-align:center;} {\Huge\cal\color {grey} {{🐬~Welcome~To\\Question~Cove!~🐬}}}\]